OpenStudy (brucebaner):

For the graph shown, select the statement that best represents the given system of equations. algebra unit 8 lesson 1 assessment question 3 6x + 4y = 2 3x + 2y = 1 A. not enough information B. coincident C. consistent and independent D. inconsistent

OpenStudy (brucebaner):

@SolomonZelman

OpenStudy (brucebaner):

@TheSmartOne

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/ljbi9z995t

TheSmartOne (thesmartone):

Look at the graph. Do they intersect are they the same line?

OpenStudy (brucebaner):

no

OpenStudy (solomonzelman):

divide the first equation by 2. they are a dang same thing

TheSmartOne (thesmartone):

^^

TheSmartOne (thesmartone):

lol

OpenStudy (brucebaner):

i think its C but im not sure

OpenStudy (solomonzelman):

one line on top of the other in 3D plane, but we are in a 2D world here..... on OS for most of the things. in a 2D world, that means? (you tell me) couple examples: \(\normalsize\color{royalblue}{ \rm x=x }\) means infinity of x solutions: \(\normalsize\color{royalblue}{ \rm 3x+4=3x +4 }\) means infinity of x solutions:

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm independent }\) means 1 solution/intersection.

TheSmartOne (thesmartone):

A dependent solution has infinitely many solutions and above :)

OpenStudy (solomonzelman):

and above?

OpenStudy (brucebaner):

so its A

TheSmartOne (thesmartone):

and above refers to what you said :) @SolomonZelman

OpenStudy (brucebaner):

so C

OpenStudy (solomonzelman):

Oh, I though you are thinking of a number greater than \(\normalsize\color{ e}{ \rm \infty }\). \(\normalsize\color{royalblue}{ \rm independent }\) means 1 solution/intersection. \(\normalsize\color{royalblue}{ \rm consistent }\) means at least solution/intersection. \(\normalsize\color{royalblue}{ \rm coincident }\) - 2 lines on top of each other.

TheSmartOne (thesmartone):

This might come in handy too: http://prntscr.com/5ojj12

OpenStudy (brucebaner):

so it is C?

OpenStudy (solomonzelman):

I provided enough info. I have a bask game w/ friends. gtg

OpenStudy (solomonzelman):

No joke... see you at some time:) and above:PPP bye!

TheSmartOne (thesmartone):

Bye!

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @SolomonZelman Oh, I though you are thinking of a number greater than \(\normalsize\color{ e}{ \rm \infty }\). \(\normalsize\color{royalblue}{ \rm independent }\) means 1 solution/intersection. \(\normalsize\color{royalblue}{ \rm consistent }\) means at least solution/intersection. \(\normalsize\color{royalblue}{ \rm coincident }\) - 2 lines on top of each other. \(\color{blue}{\text{End of Quote}}\) @brucebaner This has your answer :)

TheSmartOne (thesmartone):

Did you get it???

OpenStudy (brucebaner):

yes